Edit question what is the density and uncertainty for ideal gas equation of state if pressure is 2 atmospheres( with uncertainty 25%) and temp is 18 degrees Celsius, with uncertainty 2 degrees Celsius. Gas is nitrogen.
Answer #1

using ideal gas eqn

PM = dRT

d = PM/RT = 2x10^5 x 28/ 8315 x 291

= 2.31 kg/ m^3

dd = dP + dT = 0.25 + 2/18 = 0.25 + 0.11 = 0.36

uncertainty in density = 0.36 x 2.31 = 0.833 kg/m^3